# SATTC: Structure-Aware Label-Free Test-Time Calibration for Cross-Subject EEG-to-Image Retrieval

### 21 Mar 2026 v1

@@ Abstract
Cross-subject EEG-to-image retrieval for visual decoding is challenged by subject shift and hubness in the embedding space, which distort similarity geometry and destabilize top-$k$ rankings, making small-$k$ shortlists unreliable. We introduce SATTC (Structure-Aware Test-Time Calibration), a label-free calibration head that operates directly on the similarity matrix of frozen EEG and image encoders. SATTC combines a geometric expert—subject-adaptive whitening of EEG embeddings with an adaptive variant of Cross-domain Similarity Local Scaling (CSLS)—and a structural expert built from mutual nearest neighbors, bidirectional top-$k$ ranks, and class popularity, fused via a simple Product-of-Experts rule. On THINGS-EEG under a strict leave-one-subject-out protocol, standardized inference with cosine similarities, $\ell_{2}$-normalized embeddings, and candidate whitening already yields a strong cross-subject baseline over the original ATM retrieval setup. Building on this baseline, SATTC further improves Top-1 and Top-5 accuracy, reduces hubness and per-class imbalance, and produces more reliable small-$k$ shortlists. These gains transfer across multiple EEG encoders, supporting SATTC as an encoder-agnostic, label-free test-time calibration layer for cross-subject neural decoding. Code is available at https://github.com/QunjieHuang/SATTC-CVPR2026
@@ 1 Introduction
Early work on EEG-based visual retrieval largely relied on supervised learning, training subject-specific classifiers or regressors on labeled EEG-image pairs to predict stimulus labels or features. However, such pipelines require costly trial-level annotations, are constrained by limited data, and typically generalize poorly to unseen subjects. As real cross-subject deployments seldom provide labels for new users, unsupervised and zero-shot paradigms have emerged as natural alternatives. Several recent studies adopt encoder-centric unsupervised or zero-shot EEG-to-image paradigms that align EEG and image embeddings via contrastive learning [22, 3, 23, 13, 4, 27]. These works introduce increasingly expressive EEG encoders and report strong zero-shot recognition and retrieval performance on THINGS-EEG benchmarks [7], but their test-time evaluation typically reduces to non-standardized nearest-neighbor search over learned similarity scores in the shared space. In parallel, a separate line of EEG decoding research improves cross-subject robustness by learning subject-invariant or domain-adapted representations through adversarial inference, multi-source marginal distribution adaptation, and plug-and-play domain adaptation modules [16, 29, 5, 19]. While effective for representation learning, these methods operate at training time and have not been instantiated as generic, label-free calibration heads [25, 28] that can standardize and improve retrieval shortlists in zero-shot cross-subject EEG-to-image settings.
@@
  • We propose a structural expert that operates on the pre-CSLS similarity matrix using mutual nearest neighbors, bidirectional top-$k$ ranks, and class popularity, and fuse it with the geometric expert via a Product-of-Experts model. Experiments on the THINGS-EEG cross-subject benchmark show that our standardized inference baseline already outperforms the original ATM pipeline, and that SATTC further improves Top-1 accuracy over strong fixed-$k$ and adaptive CSLS baselines without harming Top-5 performance, while reducing hubness and per-class imbalance and yielding more reliable small-$k$ shortlists.
@@ 2.1 EEG-based Visual Decoding and Cross-Subject Generalization
EEG-based visual decoding has progressed from early coarse category recognition to fine-grained visual classification, image retrieval, and reconstruction [24, 13, 23]. A common recipe is to learn an encoder that maps short EEG segments into a visual or multimodal embedding space, followed by nearest-neighbor retrieval or regression to image features on datasets such as THINGS-EEG. Despite strong results, most systems are still evaluated in subject-specific or lightly adapted settings, assuming a separate model or labeled data for each target user, which underestimates the difficulty of deployment to unseen users without per-subject labels.
@@ 3.1 Label-Free Test-Time Calibration: Problem Setup and Overview
We consider cross-subject EEG-to-image retrieval on the THINGS-EEG dataset [9, 7], which provides multichannel EEG recordings while subjects view images from a fixed vocabulary of object categories. Let $Q$ denote EEG queries from unseen test subjects and $C$ the set of candidate image classes. A pretrained EEG encoder $f_{\text{eeg}}$ maps each query $q\in Q$ to an embedding $z_{q}\in\mathbb{R}^{d}$, and an image encoder $f_{\text{img}}$ maps each class $c\in C$ to an embedding $v_{c}\in\mathbb{R}^{d}$. After subject-adaptive whitening and normalization (Section 3.2), we obtain a pre-CSLS similarity matrix
@@ 4.1 Experimental Setup
We evaluate on the public THINGS-EEG dataset [9, 7], which provides EEG recordings from 10 subjects viewing images from 200 object categories. We adopt a leave-one-subject-out (LOSO) protocol: in each of 10 folds, one subject is held out for testing and the remaining nine are used for training and tuning.
@@ 4.3 Ablations for Label-Free Test-Time Calibration
Table 1: Cross-subject 200-way EEG-to-image retrieval results on THINGS-EEG under LOSO. Mean Top-5 and Top-1 accuracy (%) over all test folds and three random seeds for the ATM baseline, our standardized baseline (cosine + $\ell_{2}$ normalization + candidate whitening), and successive additions of SAW, candidate whitening, CSLS variants, and SATTC.
@@ 4.4 Hubness Reduction and Shortlist Quality
Table 2: Plug-and-play generalization of SATTC across EEG encoders on 200-way THINGS-EEG cross-subject retrieval. For each encoder, we report Top-5 and Top-1 accuracy (%) averaged over all LOSO folds and three random seeds, comparing our standardized inference baseline to the same encoder with SATTC applied as a label-free test-time calibration head.
@@ 4.5 Encoder-Agnostic Plug-and-Play Calibration
We next test whether SATTC is tied to a specific EEG encoder or can act as a generic calibration layer. Table 2 reports 200-way cross-subject retrieval on THINGS-EEG for four heterogeneous encoders (ATM, EEGNetV4, EEGConformer, and ShallowFBCSPNet), spanning CSP-style, compact CNN, and transformer architectures. For each encoder we compare our standardized inference baseline to the same model equipped with SATTC as a label-free test-time calibration head, without changing the backbone or using labels from test subjects.
@@ 5 Discussion and Conclusion
Limitations. SATTC is currently evaluated only on THINGS-EEG under a LOSO cross-subject retrieval protocol. Its structural expert is hand-crafted from ranks, mutual nearest neighbors, and class popularity, and the current implementation operates on precomputed similarity matrices. In practice, the SAW+CSLS branch supports calibration-once then frozen inference, whereas the structural expert and PoE are most naturally applied in a batch refinement stage when multiple unlabeled test queries are available. Extending the method to additional datasets, encoder families, and decoding regimes such as retrieval-then-generation or EEG/fMRI reconstruction [13, 4] remains an important direction for future work.
Conclusion. We cast cross-subject EEG-to-image retrieval as a label-free test-time calibration problem on similarity matrices. SATTC operates on frozen EEG and image encoders through a modular calibration head that combines subject-adaptive whitening, adaptive CSLS, and a lightweight structural prior from pre-CSLS similarities. On THINGS-EEG under strict LOSO evaluation, SATTC improves Top-1 and Top-5 over strong standardized baselines, reduces hubness and per-class imbalance, and yields more reliable small-$k$ shortlists. Together with its plug-and-play gains across multiple EEG encoders, these results suggest that similarity-space calibration is a practical route toward more robust cross-subject neural decoding.

### 25 Apr 2026 v2

@@ Abstract
Cross-subject EEG-to-image retrieval for visual decoding is challenged by subject shift and hubness in the embedding space, which distort similarity geometry and destabilize top-$k$ rankings, making small-$k$ shortlists unreliable. We introduce SATTC (Structure-Aware Test-Time Calibration), a label-free calibration head that operates directly on the similarity matrix of frozen EEG and image encoders. SATTC combines a geometric expert—subject-adaptive whitening of EEG embeddings with an adaptive variant of Cross-domain Similarity Local Scaling (CSLS)—and a structural expert built from mutual nearest neighbors, bidirectional top-$k$ ranks, and class popularity, fused via a simple Product-of-Experts rule. On THINGS-EEG2 under a strict leave-one-subject-out protocol, standardized inference with cosine similarities, $\ell_{2}$-normalized embeddings, and candidate whitening already yields a strong cross-subject baseline over the original ATM retrieval setup. Building on this baseline, SATTC further improves Top-1 and Top-5 accuracy, reduces hubness and per-class imbalance, and produces more reliable small-$k$ shortlists. These gains transfer across multiple EEG encoders, supporting SATTC as an encoder-agnostic, label-free test-time calibration layer for cross-subject neural decoding. Code is available at https://github.com/QunjieHuang/SATTC-CVPR2026
@@ 1 Introduction
Early work on EEG-based visual retrieval largely relied on supervised learning, training subject-specific classifiers or regressors on labeled EEG-image pairs to predict stimulus labels or features. However, such pipelines require costly trial-level annotations, are constrained by limited data, and typically generalize poorly to unseen subjects. As real cross-subject deployments seldom provide labels for new users, unsupervised and zero-shot paradigms have emerged as natural alternatives. Several recent studies adopt encoder-centric unsupervised or zero-shot EEG-to-image paradigms that align EEG and image embeddings via contrastive learning [22, 3, 23, 13, 4, 27]. These works introduce increasingly expressive EEG encoders and report strong zero-shot recognition and retrieval performance on THINGS-EEG2 benchmarks [7], but their test-time evaluation typically reduces to non-standardized nearest-neighbor search over learned similarity scores in the shared space. In parallel, a separate line of EEG decoding research improves cross-subject robustness by learning subject-invariant or domain-adapted representations through adversarial inference, multi-source marginal distribution adaptation, and plug-and-play domain adaptation modules [16, 29, 5, 19]. While effective for representation learning, these methods operate at training time and have not been instantiated as generic, label-free calibration heads [25, 28] that can standardize and improve retrieval shortlists in zero-shot cross-subject EEG-to-image settings.
@@
  • We propose a structural expert that operates on the pre-CSLS similarity matrix using mutual nearest neighbors, bidirectional top-$k$ ranks, and class popularity, and fuse it with the geometric expert via a Product-of-Experts model. Experiments on the THINGS-EEG2 cross-subject benchmark show that our standardized inference baseline already outperforms the original ATM pipeline, and that SATTC further improves Top-1 accuracy over strong fixed-$k$ and adaptive CSLS baselines without harming Top-5 performance, while reducing hubness and per-class imbalance and yielding more reliable small-$k$ shortlists.
@@ 2.1 EEG-based Visual Decoding and Cross-Subject Generalization
EEG-based visual decoding has progressed from early coarse category recognition to fine-grained visual classification, image retrieval, and reconstruction [24, 13, 23]. A common recipe is to learn an encoder that maps short EEG segments into a visual or multimodal embedding space, followed by nearest-neighbor retrieval or regression to image features on datasets such as THINGS-EEG2. Despite strong results, most systems are still evaluated in subject-specific or lightly adapted settings, assuming a separate model or labeled data for each target user, which underestimates the difficulty of deployment to unseen users without per-subject labels.
@@ 3.1 Label-Free Test-Time Calibration: Problem Setup and Overview
We consider cross-subject EEG-to-image retrieval on the THINGS-EEG2 dataset [9, 7], which provides multichannel EEG recordings while subjects view images from a fixed vocabulary of object categories. Let $Q$ denote EEG queries from unseen test subjects and $C$ the set of candidate image classes. A pretrained EEG encoder $f_{\text{eeg}}$ maps each query $q\in Q$ to an embedding $z_{q}\in\mathbb{R}^{d}$, and an image encoder $f_{\text{img}}$ maps each class $c\in C$ to an embedding $v_{c}\in\mathbb{R}^{d}$. After subject-adaptive whitening and normalization (Section 3.2), we obtain a pre-CSLS similarity matrix
@@ 4.1 Experimental Setup
We evaluate on the public THINGS-EEG2 dataset [9, 7], which provides EEG recordings from 10 subjects viewing images from 200 object categories. We adopt a leave-one-subject-out (LOSO) protocol: in each of 10 folds, one subject is held out for testing and the remaining nine are used for training and tuning.
@@ 4.3 Ablations for Label-Free Test-Time Calibration
Table 1: Cross-subject 200-way EEG-to-image retrieval results on THINGS-EEG2 under LOSO. Mean Top-5 and Top-1 accuracy (%) over all test folds and three random seeds for the ATM baseline, our standardized baseline (cosine + $\ell_{2}$ normalization + candidate whitening), and successive additions of SAW, candidate whitening, CSLS variants, and SATTC.
@@ 4.4 Hubness Reduction and Shortlist Quality
Table 2: Plug-and-play generalization of SATTC across EEG encoders on 200-way THINGS-EEG2 cross-subject retrieval. For each encoder, we report Top-5 and Top-1 accuracy (%) averaged over all LOSO folds and three random seeds, comparing our standardized inference baseline to the same encoder with SATTC applied as a label-free test-time calibration head.
@@ 4.5 Encoder-Agnostic Plug-and-Play Calibration
We next test whether SATTC is tied to a specific EEG encoder or can act as a generic calibration layer. Table 2 reports 200-way cross-subject retrieval on THINGS-EEG2 for four heterogeneous encoders (ATM, EEGNetV4, EEGConformer, and ShallowFBCSPNet), spanning CSP-style, compact CNN, and transformer architectures. For each encoder we compare our standardized inference baseline to the same model equipped with SATTC as a label-free test-time calibration head, without changing the backbone or using labels from test subjects.
@@ 5 Discussion and Conclusion
Limitations. SATTC is currently evaluated only on THINGS-EEG2 under a LOSO cross-subject retrieval protocol. Its structural expert is hand-crafted from ranks, mutual nearest neighbors, and class popularity, and the current implementation operates on precomputed similarity matrices. In practice, the SAW+CSLS branch supports calibration-once then frozen inference, whereas the structural expert and PoE are most naturally applied in a batch refinement stage when multiple unlabeled test queries are available. Extending the method to additional datasets, encoder families, and decoding regimes such as retrieval-then-generation or EEG/fMRI reconstruction [13, 4] remains an important direction for future work.
Conclusion. We cast cross-subject EEG-to-image retrieval as a label-free test-time calibration problem on similarity matrices. SATTC operates on frozen EEG and image encoders through a modular calibration head that combines subject-adaptive whitening, adaptive CSLS, and a lightweight structural prior from pre-CSLS similarities. On THINGS-EEG2 under strict LOSO evaluation, SATTC improves Top-1 and Top-5 over strong standardized baselines, reduces hubness and per-class imbalance, and yields more reliable small-$k$ shortlists. Together with its plug-and-play gains across multiple EEG encoders, these results suggest that similarity-space calibration is a practical route toward more robust cross-subject neural decoding.